Jefferson Lab
JLab Web Server 
Jefferson Lab Home Search Contact JLab

FEL Upgrade Project Weekly Brief
January 20-24, 2003

We got another productive week of installation, qualification and commissioning activities on the FEL upgrade
behind us with our eye on the start up of significant beam commissioning activities by the middle of February.  We
completed our last SRF commissioning tasks on the injector cryounit, recycled the gun into its second bakeout
cycle, prepared for installation of the last three magnets and chambers in the first recirculation arc, and completed
wiring in the high reflector optical cavity.

Commissioning Activities:
1.) The last SRF commissioning task was completed this week. The second SRF cavity (#4) in the injector cryounit
     was fully qualified to a gradient (12.4MV) above what is needed for its previous operation limit (11MV).
2.) The vacuum chambers are qualified and ready for installation of three remaining magnets (1 GQ and 2GX's)
    which is planned to start on next Wednesday and complete by Feb. 5th.  This will complete the first recirculation arc
3) The third vacuum chamber in the optical chicane was installed this week. The last chamber is planned to be
    installed by Feb. 4 or 5th. This would then complete the last major installation activity in the back leg.
4) The wiring for the high reflector optical cavity assembly was completed this week and this task is nearly
    complete with check-out.
5) The GaAs cathode was replaced in the gun last weekend and the gun vacuum chamber was rebaked to obtain
    the necessary low vacuum conditions to re start HV conditioning activities when we ramp down the bakeout next

Project management reports for the months of October and November were distributed to the DOE and DOD
program offices.  With all the special reports that were needed after the November review, we fell behind in issuing
our standard reports.  We hope to get the December report out next week to get back on schedule.
We sent back to the panel chair minor comments and corrections to the draft November project review panel
report which was distributed late last month.  We thank the review panel members and the Chair for their diligent
effort which was evident in the preparation of the report.
On Jan. 24, Jefferson Lab hosted a workshop on the potential for x-ray lithography to reach the 15 nm regime.
Workshop participants included representatives from IBM, Lucent, BAE, several stepper manufacturers, several
universities with interest in lithography, NRL, and DARPA.

WBS 4 (Injector):
A new wafer was installed into the gun and a vacuum bakeout started.  The bake should be done on Monday when
high voltage conditioning will be started.

WBS 6 (RF):
Quarter - By lowering the broadband gain of the RF Control Module, cavity 4 was more stable and able to achieve
its drive high gradient.  Auto-Tracking will be tried again this evening.  Auto-Tune may take some additional

Zone 4 - In order to achieve the high gradients for this zone, the cathode voltage will have to be raised above our
standard 11.6 kV.  They should be able to operate to 14 kV.  Even with the cavities carefully tuned, several of
them cavities were arcing at much lower gradients than their expected drive high values.  The cathode voltage will
be increased this evening when testing continues.

WBS 8 (Instrumentation):
All 4-Channel BPMs in the 0F, 1F and 2F regions have been connected and tested.  EPICS readouts are in the
process of being updated to reflect the correct readings for each monitor.  The remainder of the 4 channel BPM
leads (LMR 200) for the 4F and 5F regions have been placed
on order.

All magnets are installed in the LINAC region to the tune-up dump.  Power connections will be completed and
tested early next week.  All labels are being updated as well.  The DC power line to 2F05 and the DY (first arc
bend) was completed from the 66kW box supply.  A high level  check out can be performed early next week.  DC
power for the GX's, GQ's and optical chicane  for the 1st Arc string are being routed.  High level check out of the
DC power for the 1st Arc String can take begin when MGX2F06 and MGQ2F07 are installed.  Core lead wiring
of the Sextapole magnets are in progress.

Thanks to EECAD for their help in the rapid completion of the following drawings:  30Hz Signal Distribution Box,
IPPS Controller PCB Schematic, LSS User Lab Shutter Control Box.

The HVPS permits (OFF & NOT ON) from the injector area have been terminated to the 24V to Fiber converter
card in iocfel10.  After which, they provide the DLPC with the ability to override the MPS when the HVPS is not
on. (This allows the drive laser to be run independent of the MPS
when we are not in BEAM PERMIT).  The HVPS Interlock analog I/O module was received from the vendor,
populated, installed and check out is in progress.  The calibration to full current will be done tonight (1/24).

The second Stepper Motor chassis for the Output Coupler has been completed.
 * Upgrade and clean-up continues on the video system with primary focus on the video rack FL02B09.  The User
Lab Area Monitors on the wall in the control room are all in full operation. The TV in the clean room has been
restored into the system on channel 9.  The Injector Brewster window and Cathode Camera are connected into the
system now as well as another viewer line (ITV0F06).  Additionally the 2nd video switcher chassis is almost
reinstated pending the remake of an old cable assembly. This will increase the number of functioning channels in the
system from 64 to 128. * The software controlling the variable attenuator has been debugged and is now working
correctly.  It will come up properly after a reboot so long as the ioc (iocfel10) is brought up properly. * Cameras
have been put on the strip-chart and ion pumps being used in the gun bake-out.  These cameras have been plugged
into our online video system and can be viewed at:

Extension 2 for the Interim ODH Ratings for the FEL, TSOP A-02-009-T-R2 has been approved and becomes
effective tomorrow, January 25, 2003.  The ODH Risk Assessment for the FEL has been completed.  The FEL
Facility will be ODH 0 during normal operations when the following conditions are met:  The return valve for the
helium relief valves is locked open and an administrative lockout tag applied and all penetrations from the vault to
the second floor have been sealed.  These items should be completed by the end of next week and an SOP will be
issued as required by the assessment.

WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in the First Arc.
•  The power supply and cabling to this magnet (and the alternate DW spectrometer magnet at the end of the
Linac)  was completed.
Diversion of the beam after the Wiggler.
•  The vacuum piping beyond this magnet started taking shape.
Optical Chicane Dipoles (GW)
•  No progress.
Injector Dipoles (DU/DV)
•  No action
Arc 180 Degree Dipoles (GY)
•  Don Bullard continued putting together the semicircular-tracked magnet measurement apparatus between bouts
of work on the Gun.
• Wang is going to pot the first coil of the second GY next Tuesday.  They predict shipment in three weeks.
Arc Bend & Reverse Bend Dipoles (GQ, GX)
•  David Douglas decided that the straight line integrals we rely on are suspect.  The integrals are extracted from the
    data on these two series of 43 degree bend angle magnets after adding face shim and field clamp shims that raise
    or lower the local field to obtain the required integral within 2 parts in 10,000.  At the large bend angle, entering
    beam could be going through a high end field compensation and then traverse a high central field zone and thus be
    subject of excessive BdL.  Adjacent beam could be going through a low field zone at the center of the magnet at the
    entrance and exit and a lower field zone at the side of the pole tip and thus be exposed to deficient BdL.  As an
    alternate analysis, David will analyze the magnets along the beam paths by interpolating the data matrix.  This will
    take time.  To get on with installation, he is releasing the magnets for installation (without sign off) after we have a
    complete set of measurements for analysis.  The magnets also have to show they have the necessary (not
    necessarily sufficient) bulk field, field integral and field integral slope.
•  A misunderstanding of the value of the required field integrals (DIMAD cord vs. half sagitta cord) and some
    drawing misinterpretation were cleared up during our analysis of the data.  The result contrary to our original
    findings reported thus far.  The GQ magnets are short by about 7 mm and the GX magnets are just right.

    Fortunately, we have shims available for the GQs that lengthen the field by about 5.5 mm (They were to be used
    for the 210 MeV/c case).  At week’s end we added an additional 0.75 mm to each shim stack and found the
    second GQ magnet is well within specification at three parts in 10,000 of the required integral.  With this predictability,
    David is willing to allow the same shim addition to the magnet already in the enclosure without additional
    measurements.  Behavior at 210 MeV/c will be done on the second set of two magnets after the shim package
    that adds even more shim is developed.

    The GX magnets have the required field integral along the half sagitta line but their field integral gradient
    transverse to the beam is out of specification.  Measurements next week will focus on altering the angle of the field clamps
    to dial out this droop. Again, later analysis along the beam path will be the final determiner of official sign-off by
    David Douglas.

• Wang NMR potted all four GQ coils and is predicting shipment of both complete magnets at the end of next
Trim Quad (QT)
•  The trim coils to create the GC vertical corrector function within the four trim quads were received from Milhous
    Control.  The trim quads will be taken apart and converted to the combination magnets by the Jlab Shop.
•  The remaining Trim Quads were mounted on their stands in the enclosure
Sextupole (SF)
•  We are setting up a method to wire up the magnets as they are on their stands in the enclosure.
Octupole (OT)
•  The procurement process continued.
Beam Line and Vacuum
•  Our shop continued working on the last chamber in Optical Chicane.  They forecast completion on the 4thth of
•  The magnets of the Optical Chicane were wired up to the Arc 1 power supply.
•  The two 180o  Bend chambers continue to await the receipt of the diagnostic windows from Ceramaseal.  The
    latest projected delivery date is early February.
• The Vacuum Group connected girders in the regions around the optical klystron in their drive to button up the
•  All the correctors and skew quads in the Linac were installed and most are wired up.

WBS 11 (Optics):
This week we attempted to connect our hardware to the controls systems.  We had some glitches; cable
terminations from upstairs were different than our prints; we had some miswirings of our own, and one LVDT
became partially disassembled and would not read back correctly.  There will be a report in WBS 8 (I&C) on their
progress in working some controls issues.  In summary, we are encountering and working through standard
commissioning issues.  We are still waiting to receive components to bypass cooling water for deformable mirror
assemblies we aren't ready to install.  This will create about a week slip in our schedule; thus the HR assembly will
be put under vacuum and we continue commissioning next week.  Fabricated parts for the diagnostic viewer/laser
safety shutter were received today.  After an initial "dirty" assembly, we will have the parts UHV cleaned and install
them.  All parts for the insertable mirror are also here, except for the Dicronited setscrews.  They will arrive nearly
next week.  The optics and laser for the HR ultraviewer were installed on their optics rail and aligned.  They will be
installed after the dipole chicane chamber for that region is installed.

Other commissioning activities:
We checked out repeatability and operation of the THz diagnostic mirror.  We will install a pyroelectric detector on
it during the initial injector testing.  Work continues on mounting and aligning the more sophisticated fiber optics for
this diagnostic.  A new, servo-controlled translation stage for repeatably yet gently inserting a chopper has been
tested successfully.  This can be used to replace one of the EO modulators if we wish more laser throughput, as
well as a somewhat simpler EO optimization routine.  Drive laser optical transport system alignment is scheduled to
take place tomorrow (1/25).